OpenStudy (anonymous):

The triangle shown is a right triangle. Create the equation to be used to find the missing lengths. ( Enter the smaller leg of the triangle first) do not solve the equation. The hypotenuse is x, the short side is 4 and the long side is 7. I answered x^2= 4^2+ 7^2

OpenStudy (anonymous):

 equation for finding x would be by the Pythagorean theorem since its a right triangle. a^2+b^2=c^2 7.1^2+4^2=hypotenuse^2 then find the root of the hypotenuse

OpenStudy (anonymous):

Thanks, idk why it is counting it wrong then